# A Dissipative Dark Cosmology: From Early Matter Dominance to Delayed Compact Objects

### v1

@@ Abstract
We demonstrate a novel mechanism for producing dark compact objects and black holes through a dark sector, where all the dark matter can be dissipative. Heavy dark sector particles with masses above $10^{4}$ GeV can come to dominate the Universe and yield an early matter-dominated era before Big Bang Nucleosynthesis (BBN). Density perturbations in this epoch can grow and collapse into tiny dark matter halos, which cool via self interactions. The typical halo size is set by the Hubble length once perturbations begin growing, offering a straightforward prediction of the halo size and evolution depending on ones choice of dark matter model. Once these primordial halos have formed, a thermal phase transition can then shift the Universe back into radiation domination and standard cosmology. These halos can continue to collapse after BBN, resulting in the late-time formation of fragmented dark compact objects and sub-solar mass primordial black holes. We find that these compact objects can constitute a sizable fraction of all of dark matter. The resulting fragments can have masses between $10^{20}$ g to $10^{32}$ g, with radii ranging from $10^{-2}$ m to $10^{5}$ m, while the black holes can have masses between $10^{8}$ g to $10^{34}$ g. Furthermore, a unique feature of this model is the late-time formation of black holes which can evaporate away today. We compare where these objects lie with respect to current primordial black hole and MACHO constraints.
@@ I Introduction
Cold, collisionless dark matter (DM) works well to explain structure formation over a wide range of scales. However, dark matter’s non-gravitational interactions, especially its possible self interactions, are poorly understood. Given the complexity and variety of forces within the Standard Model (SM), it is worth considering similar interactions within the dark sector. In fact, DM self interactions have previously been invoked to explain potential problems in cosmology, such as issues with small scale structure formation [1, 2, 3, 4, 5] and the Hubble tension [6]. More generally, related studies have explored how such a dissipative dark sector could produce collapsed structures like dark stars and black holes, though typically this dissipative sector is assumed to be only a subdominant component of the total DM budget [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
Recently, several works have studied dark structure formation prior to big bang nucleosynthesis (BBN). While some have explored the behavior of dissipative dark sectors assuming radiation domination at these early times [19, 20, 21], others have indicated that collisionless dark matter can form structures in an early matter-dominated era (EMDE) [22, 23, 24, 25]. While EMDEs have been getting more attention lately, they have a long history within cosmology theory, reaching back to early work on primordial black holes (PBHs) [26, 27]. PBHs can come to dominate the Universe and evaporate away before BBN [28, 29] or form from structure collapse during an EMDE and persist to late times [30, 31, 32].
@@ II Dark sector overdensity and growth
Following a standard inflationary epoch and a subsequent period of reheating, we assume the SM and the dark sector are in thermal equilibrium in a relativistic bath of particles. Once the temperature of the particle bath drops below the dark matter mass, its abundance becomes Boltzmann-suppressed. In the presence of a large enough initial asymmetry (e.g. an $\mathcal{O}(1)$ asymmetry as expected in Affleck-Dine models [40, 38]), then the comoving number density will rapidly freeze-out, and proceed to become the dominant energy density of the Universe as the relativistic plasma is redshifted away, leading to an era of early matter domination.
The growth of perturbations and the collapse of overdensities proceed in this EMDE as they would in an ordinary matter dominated era. We assume that at the start of the EMDE, the initial density perturbations, $\delta=(\rho-\bar{\rho})/\bar{\rho}$, are of order $\delta_{0}\sim 10^{-5}$, similar to typical primordial perturbations observed at late times [41]. We note that one may choose different, potentially larger, initial density perturbations for these small modes entering the horizon at such early times as in some treatments of EMDEs [22, 30]. Perturbations grow linearly with the scale factor $a$, until turnaround, when $\delta\sim 1$, at which point the perturbations begin to collapse, and virialize independently of the background expansion of the Universe. A cartoon depiction of the full evolution of dark electron structures from density perturbation to final collapsed object can be found in Fig. 1.
@@
$\frac{\textrm{d}\sigma}{\textrm{d}\Omega}=\frac{m_{X}^{2}\alpha_{D}^{2}(4p^{4}(1+3\cos^{2}\theta)+4p^{2}m_{\gamma_{D}}^{2}+m_{\gamma_{D}}^{4})}{(4p^{4}\sin^{2}\theta+4p^{2}m_{\gamma_{D}}^{2}+m_{\gamma_{D}}^{4})^{2}}\,.$ (6)
@@ III Dark sector collapse and final masses
The collapse of primordial dark electron clouds can be separated into three stages. First, the $X$ overdensities free-fall under their self-gravity at timescales of $t_{\mathrm{collapse}}=t_{\mathrm{ff}}=(16\pi G\rho_{X})^{-1/2}$ so that the collapse trajectory is $d\log T_{X}/d\log n_{X}\simeq 2/3$. The overdensities will free-fall into virialized halos with radii about half of that of the overdensity at turnaround [43]. During free fall (as long as the gas is collisional) the gas will heat up rapidly due to shock heating. Exchanging gravitational potential energy for thermal energy in the gas drives the temperature of the halo to about $T_{\mathrm{vir}}\sim m_{X}\delta_{0}/3$. If $T_{\mathrm{vir}}>m_{\gamma_{D}}$, bremsstrahlung is efficient and the gas then undergoes a nearly-virialized contraction phase with $t_{\mathrm{collapse}}=t_{\mathrm{cool}}/6$. The gas collapses along a constant Jeans mass contour $m_{J}\sim M_{H}$ with $d\log T_{X}/d\log n_{X}\simeq 1/3$. From here, the halo will continue contracting until either collapsing into a black hole or fragmenting into a collection of smaller pressure supported objects. When $t_{\mathrm{cool}}>t_{\mathrm{ff}}$ for the entire collapse process, the halo collapses directly into a black hole. Otherwise, if the temperatures and densities of the halos are large enough to make cooling efficient–as fast as collapse ($i.e.$ $t_{\mathrm{collapse}}=t_{\mathrm{cool}}\simeq t_{\mathrm{ff}}$), the gas fragments into smaller clumps. If cooling is very inefficient such that $t_{\mathrm{cool}}$ is larger than the age of the Universe, then the primordial halo today will be a pressure supported object with about the same temperature and density it had when it first virialized.
@@ IV Pre-BBN phase transition
Thermal inflation will end when the temperature of thermal bath particles drops below the temperature at which the second-order correction terms in Eq. (14) are approximately equal, so that $T_{\mathrm{PT}}\approx m/\sqrt{2\lambda_{s}}$. Once thermal inflation ends, we assume the decay of the scalar $\phi$ with potential energy $V_{0}$ into SM radiation happens quickly, and so the energy density of radiation after the decay is given by
@@
where in the final inequality we have taken $T_{\mathrm{mr}}\simeq 0.8$ eV. We have also used the fact that the background density of $X$ particles following the phase transition is given by $\bar{\rho}_{X,\mathrm{RH}}=\pi^{2}g^{*}_{\mathrm{PT}}T_{\mathrm{PT}}^{3}T_{\mathrm{FO}}/30$, where again $T_{\mathrm{FO}}\approx m_{X}/10$ is the temperature at which the $X$ particles froze out and became non-relativistic.
@@ V.1 Dark compact objects
These DarkCOs are very compact objects and can thus be treated as point sources – making them susceptible to microlensing constraints. Indeed, we can apply existing constraints on MACHOs [52] based on data from microlensing surveys such as Subaru/HSC [54], MACHO [55], EROS-2 [56], OGLE [57], and Kepler [58], to the DarkCOs predicted in this work. The overlaid white bars in the color bar of Fig. 4 indicate ruled-out DarkCO mass ranges assuming a fraction $f$ of dark matter is in these compact objects. We have considered $f=0.17$ and $f=0.05$ corresponding to the assumption that dark electrons comprise all of the dark matter and that $17\%$ or $5\%$ of the dark electrons have formed compact structures. These fractions correspond to scenarios in which thermal inflation begins after all density perturbations with an initial overdensity at least $1\sigma$ and $2\sigma$ above the background density have had time to collapse. For a fraction of $f=0.17$, only lower mass DarkCOs (very top right of Fig. 4) are permitted due to the strong constraint coming from Subaru/HSC, OGLE, and EROS-2. However, for a lower value of $f=0.05$, the constraint from OGLE is weaker, and hence an additional window opens up between $10^{27}$ g (coming from Subaru/HSC) and $10^{29}$ g (coming from EROS-2).
@@ V.2 Black Holes
Meanwhile, Fig. 5 shows a zoomed-in version of the predicted black holes in Fig. 3, with collapse timescales at virialization overlaid on top. Similar to Fig. 4, we have indicated allowed PBH masses assuming a PBH fraction of $f=0.17$ and $f=0.05$. The lower mass exclusion band arises from evaporation constraints (see [53] for a thorough review), while the same microlensing constraints discussed for Fig. 4 apply for higher masses. PBHs with masses $M\lesssim 10^{10}$ g are unconstrained because they decay too early to impact BBN. The allowed asteroid-mass window between the two constrained regions for both values of $f$ arise from the gap between evaporation constraints and microlensing constraints.
The collapse timescales increase as the dark photon masses decrease, primarily since the initial density of the clump decreases as a function of the dark photon mass. Interestingly, objects with cooling times longer than the age of the Universe at BBN imply late time formation of PBHs. Late-forming PBHs have been studied in previous works, though these relied on a different PBH production mechanism [59, 60].
@@
Decreasing $\alpha_{D}$ makes cooling less efficient, and delays the dark electron halo collapse into a PBH. As shown in the left panel of Fig. 6, for some parts of dark electron parameter space, reducing $\alpha_{D}$ to $10^{-7}$ leads to the formation of low mass black holes which take longer to collapse than to decay. As the evaporation of these late forming PBHs would happen at different times than usually expected for PBHs, evaporation constraints are altered. The majority of the parameter space in the lower right corner of the left panel of Fig. 6 is unconstrained since these dark halos would not have collapsed into black holes yet. The modified constraints on late forming decaying black holes are shown in the right panel of Fig. 6. The constraints presented here come from modifying existing evaporating black holes limits based on observations of primordial elemental abundances from BBN [61], of CMB anisotropies [62], and from gamma-ray emission at late times [63]. Here we derive these constraints by making the simplifying assumption that the majority of energy emitted by these black holes is released at a time $t_{\mathrm{cool}}+t_{\mathrm{decay}}$, where $t_{\mathrm{decay}}=5120G^{2}M^{3}$ is the decay time for black hole. The dependence of the constraint on the energy injection time is evident in Fig. 6. The constraint contours follow the black hole mass contours on the left side of the plot where $t_{\mathrm{cool}}<t_{\mathrm{decay}}$, and the decay time sets the constraint on the black hole abundance. On the right side of the right panel in Fig. 6 where $t_{\mathrm{cool}}>t_{\mathrm{decay}}$ the constraint contours largely follow cooling time contours. As in the left panel of Fig 6, there are no evaporation constraints in the bottom right corner of right panel in Fig. 6 as black holes have yet to form in this part of parameter space. A more accurate estimate of these constraints would require fully simulating the unique radiation history of these delayed black holes and its impact on the primordial elemental abundances, the CMB, and on the observed gamma ray-spectrum, instead of simply assuming the energy is emitted instantaneously at $t_{\mathrm{cool}}+t_{\mathrm{decay}}$. We leave such careful treatments to future work.
[…]
@@
Finally, in addition to the GW emission from the exotic cosmology, binary mergers of both the compact objects and black holes produced in this scenario would give rise to GW emissions [68, 69, 70, 71, 72]. Furthermore, the evaporation of PBHs can also affect the background of stochastic GWs [29, 73]. We leave these considerations for future studies.

### v2

# Dissipative Dark Cosmology: From Early Matter Dominance to Delayed Compact Objects

@@ Abstract
We demonstrate a novel mechanism for producing dark compact objects and black holes through a dark sector, where all the dark matter can be dissipative. Heavy dark sector particles with masses above $10^{4}$ GeV can come to dominate the Universe and yield an early matter-dominated era before Big Bang Nucleosynthesis (BBN). Density perturbations in this epoch can grow and collapse into tiny dark matter halos, which cool via self interactions. The typical halo size is set by the Hubble length once perturbations begin growing, offering a straightforward prediction of the halo size and evolution depending on ones choice of dark matter model. Once these primordial halos have formed, a thermal phase transition can then shift the Universe back into radiation domination and standard cosmology. These halos can continue to collapse after BBN, resulting in the late-time formation of fragmented dark compact objects and sub-solar mass primordial black holes. We find that these compact objects can constitute a sizable fraction of all of dark matter. The resulting fragments can have masses between $10^{20}$ g to $10^{32}$ g, with radii ranging from $10^{-2}$ m to $10^{5}$ m, while the black holes can have masses between $10^{8}$ g to $10^{34}$ g. Furthermore, a unique feature of this model is the late-time formation of black holes which can evaporate today. We compare where these objects lie with respect to current primordial black hole and and massive (astrophysical) compact halo object constraints.
@@ I Introduction
Cold, collisionless dark matter (DM) works well to explain structure formation over a wide range of scales. However, dark matter’s non-gravitational interactions, especially its possible self interactions, are poorly understood. Given the complexity and variety of forces within the Standard Model (SM), it is worth considering similar interactions within the dark sector. In fact, DM self interactions have previously been invoked to explain potential problems in cosmology, such as issues with small scale structure formation [1, 2, 3, 4, 5] and the Hubble tension [6]. More generally, related studies have explored how such a dissipative dark sector could produce collapsed structures like dark stars and black holes, though typically this dissipative sector is assumed to be only a subdominant component of the total DM budget [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
Recently, several works have studied dark structure formation prior to big bang nucleosynthesis (BBN). While some have explored the behavior of dissipative dark sectors assuming radiation domination at these early times [20, 21, 22, 23, 24], others have indicated that collisionless dark matter can form structures in an early matter-dominated era (EMDE) [25, 26, 27, 28]. While EMDEs have been getting more attention lately, they have a long history within cosmology theory, reaching back to early work on primordial black holes (PBHs) [29, 30]. PBHs can come to dominate the Universe and evaporate away before BBN [31, 32] or form from structure collapse during an EMDE and persist to late times [33, 34, 35].
@@ II Dark sector overdensity and growth
Following a standard inflationary epoch and a subsequent period of reheating, we assume the SM and the dark sector are in thermal equilibrium in a relativistic bath of particles. Once the temperature of the particle bath drops below the dark matter mass, its abundance becomes Boltzmann-suppressed. In the presence of a large enough initial asymmetry (e.g. an $\mathcal{O}(1)$ asymmetry as expected in Affleck-Dine models [43, 41]), then the comoving number density will rapidly freeze-out, and proceed to become the dominant energy density of the Universe as the relativistic plasma is redshifted away, leading to an era of early matter domination. We note that in this work the dark electrons will also come to constitute most of the dark matter, since although $\gamma_{D}$ is also a potential dark matter component, we will only consider $m_{\gamma_{D}}\ll m_{X}$, meaning the dark photon will redshift like matter at a lower temperature than the dark electron, from which we expect the relative dark photon///electron late-time abundance to be $\sim m_{\gamma_{D}}/m_{X}$.
The growth of perturbations and the collapse of overdensities proceed in an $X$-dominated EMDE as they would in an ordinary SM matter dominated era. We assume that at the start of the EMDE, the initial density perturbations, $\delta=(\rho-\bar{\rho})/\bar{\rho}$, are of order $\delta_{0}\sim 10^{-5}$, similar to typical primordial perturbations observed at late times [44]. We note that one may choose different, potentially larger, initial density perturbations for these small modes entering the horizon at such early times as in some treatments of EMDEs [25, 33]. Perturbations grow linearly with the scale factor $a$, until turnaround, when $\delta\sim 1$, at which point the perturbations begin to collapse, and virialize independently of the background expansion of the Universe. A cartoon depiction of the full evolution of dark electron structures from density perturbation to final collapsed object can be found in Fig. 1.
@@
$\frac{d\sigma}{d\Omega}=\frac{m_{X}^{2}\alpha_{D}^{2}(4p^{4}(1+3\cos^{2}\theta)+4p^{2}m_{\gamma_{D}}^{2}+m_{\gamma_{D}}^{4})}{(4p^{4}\sin^{2}\theta+4p^{2}m_{\gamma_{D}}^{2}+m_{\gamma_{D}}^{4})^{2}}\,.$ (6)
@@ III Dark sector collapse and final masses
The collapse of primordial dark electron clouds can be separated into three stages. First, the $X$ overdensities will free-fall into virialized halos with radii about half of that of the overdensity at turnaround [46]. During free fall (as long as the gas is collisional) the gas will heat up rapidly due to shock heating. Exchanging gravitational potential energy for thermal energy in the gas drives the temperature of the halo to about $T_{\mathrm{vir}}\sim m_{X}\delta_{0}/3$. If $T_{\mathrm{vir}}>m_{\gamma_{D}}$, bremsstrahlung is efficient and the gas then undergoes a nearly-virialized contraction phase with $t_{\mathrm{collapse}}=t_{\mathrm{cool}}/6$. The gas collapses along a constant Jeans mass contour $m_{J}\sim M_{H}$ with $d\log T_{X}/d\log n_{X}\simeq 1/3$. From here, the halo will continue contracting until either collapsing into a black hole or fragmenting into a collection of smaller pressure supported objects. When $t_{\mathrm{cool}}>t_{\mathrm{ff}}$ for the entire collapse process, the halo collapses directly into a black hole. Otherwise, if the temperatures and densities of the halos are large enough to make cooling efficient–as fast as collapse ($i.e.$ $t_{\mathrm{collapse}}=t_{\mathrm{cool}}\simeq t_{\mathrm{ff}}$), the gas fragments into smaller clumps. If cooling is very inefficient such that $t_{\mathrm{cool}}$ is larger than the age of the Universe, then the primordial halo today will be a pressure supported object with about the same temperature and density it had when it first virialized.
@@ IV Pre-BBN phase transition
Thermal inflation will end when the temperature of thermal bath particles drops below the temperature at which the second-order correction terms in Eq. (14) are approximately equal, so that $T_{\mathrm{PT}}\approx m_{\phi}/\sqrt{2\lambda_{s}}$. Once thermal inflation ends, we assume the decay of the scalar $\phi$ with potential energy $V_{0}$ into SM radiation happens quickly, and so the energy density of radiation after the decay is given by
@@
where in the final inequality we have taken $T_{\mathrm{mr}}\simeq 0.8$ eV. We have also used the fact that the background density of $X$ particles following the phase transition is given by $\bar{\rho}_{X,\mathrm{RH}}=\pi^{2}g^{*}_{\mathrm{PT}}T_{\mathrm{PT}}^{3}T_{\mathrm{FO}}/30$, where again $T_{\mathrm{FO}}\approx m_{X}/10$ is the temperature at which the $X$ particles froze out and became non-relativistic. Note that the condition in Eq. (18) determines the relic abundance of the dark electrons. If we take a strict equality in Eq. (18), then this would correspond with a scenario where the dark electrons constitute all of dark matter, as the contribution of the dark photons in the overall dark matter density are negligible.
@@ V.1 Dark compact objects
These DarkCOs are very compact objects and can thus be treated as point sources – making them susceptible to microlensing constraints. Indeed, we can apply existing constraints on MACHOs [55] based on data from microlensing surveys such as Subaru/HSC [57], MACHO [58], EROS-2 [59], OGLE-III+IV [60, 61, 62], and Kepler [63], to the DarkCOs predicted in this work. The overlaid white bars in the color bar of Fig. 4 indicate ruled-out DarkCO mass ranges assuming a fraction $f$ of dark matter is in these compact objects. We have considered $f=0.17$ and $f=0.05$ corresponding to the assumption that dark electrons comprise all of the dark matter and that $17\%$ or $5\%$ of the dark electrons have formed compact structures. These fractions correspond to scenarios in which thermal inflation begins after all density perturbations with an initial overdensity at least $1\sigma$ and $2\sigma$ above the background density have had time to collapse. For a fraction of $f=0.17$, only lower mass DarkCOs (very top right of Fig. 4) are permitted due to the strong constraint coming from Subaru/HSC, OGLE-III+IV, and EROS-2. However, for a lower value of $f=0.05$, the constraint from EROS-2 is weaker, and hence an additional window opens up between $10^{27}$ g (coming from Subaru/HSC [57]) and $10^{29}$ g (from the 20 year dataset of OGLE-III+IV [61, 62]).
We have treated the DarkCOs as point sources due to their high density and compactness, and checked that their characteristic radii are smaller than their Einstein radii. However, for larger extended dark objects, their non-trivial structure can result in weakened microlensing constraints [64, 65]. Indeed, Ref. [66] showed that extended compact objects in our DarkCO mass range can make up all of the dark matter, if they have large enough radii. While the DarkCOs presented in the parameter space we consider are too small in mass to be found in current microlensing surveys, it is possible that mergers of the DarkCOs or accretion may further increase their sizes. We leave a more careful treatment of the mass and radius distributions of merged DarkCOs and their resulting constraints to future work. CMB measurements can also be used to constrain the abundance of DarkCOs [67], but only at masses somewhat above the mass range of the model parameters we have considered.
@@ V.2 Black Holes
Meanwhile, Fig. 5 shows a zoomed-in version of the predicted black hole parameter space in Fig. 3, with collapse timescales at virialization overlaid on top. Similar to Fig. 4, we have indicated allowed PBH masses assuming a PBH fraction of $f=0.17$ and $f=0.05$. The lower mass exclusion band arises from evaporation constraints (see [56] for a thorough review), while the same microlensing constraints discussed for Fig. 4 apply for higher masses. PBHs with masses $M\lesssim 10^{10}$ g are unconstrained because they decay too early to impact BBN. The allowed asteroid-mass window between the two constrained regions for both values of $f$ arise from the gap between evaporation constraints and microlensing constraints. Recently, Ref. [68] reexamined CMB constraints on primordial black holes to set strong limits on the PBH abundance, but only for masses somewhat above the largest masses we consider.
The collapse timescales increase as the dark photon masses decrease, primarily since the initial density of the clump decreases as a function of the dark photon mass. Interestingly, objects with cooling times comparable to the age of the Universe at BBN imply the late-time formation of PBHs, with some parameter space having black holes forming today or in the future. Late-forming PBHs have been studied in previous works, though these relied on a different PBH production mechanism [69, 70, 71].
@@
However, decreasing $\alpha_{D}$ modifies the existing PBH bounds in multiple ways. First, it makes cooling less efficient, causing the halos to collapse into PBHs at later times. As shown in the left panel of Fig. 6, for some regions of the dark electron parameter space, reducing $\alpha_{D}$ to $10^{-7}$ leads to the formation of low mass black holes which take longer to collapse than to decay. As the evaporation of these late forming PBHs would happen at different times than usually expected for PBHs, evaporation constraints are altered. The majority of the parameter space in the lower right corner of the left panel of Fig. 6 is unconstrained since these dark halos would not have collapsed into black holes yet.
Second, when $\alpha_{D}$ is reduced to $10^{-7}$, the PBHs generated from the heaviest dark electrons will be near-extremal (as defined below) when they form. This causes them to have a cooler horizon temperature, radiate a smaller fraction of their mass, and emit radiation more slowly than uncharged black holes of the same mass. The modified constraints on late forming and near-extremal decaying black holes are shown in the right panel of Fig. 6. The constraints presented here come from modifying existing evaporating black holes limits based on observations of primordial elemental abundances from BBN [72], of CMB anisotropies [73], and from gamma-ray emission at late times [74].
When describing the black holes as near-extremal, we are referencing the fact that black holes formed when a dark electron cloud collapses will retain the charge of the dark electrons that formed it. Such charged black holes cannot evaporate away to arbitrarily small masses as is often predicted for Schwarzschild black holes, but will instead evaporate down to an extremal mass, dictated by the the black hole’s starting mass, $M_{\mathrm{BH}}$, as well as $m_{X}$, $m_{\gamma_{D}}$ and $\alpha_{D}$. If the mass of the black hole is far from its extremal limit, the event horizon temperature and radius will resemble those of a Schwarzschild black hole. When the mass comes within an order of magnitude of the extremal limit, becoming “near-extremal”, its behavior will deviate significantly from that of a Schwarzschild black hole. As it approaches its extremal mass, the event horizon radius will stop shrinking and the horizon temperature will approach zero. See Appendix A for details. Note that when $\alpha_{D}=0.1$, none of the black holes predicted in this work will be near-extremal, leaving existing black hole decay constraints unchanged.
The constraints on late forming and near-extremal black holes when $\alpha_{D}=10^{-7}$ from BBN [72], of CMB anisotropies [73], and from gamma-ray emission at late times [74] are presented in the right panel of Fig. 6. We make the simplifying assumption that the black hole energy is deposited mostly at a time $t\propto t_{\text{cool}}+t_{\text{decay}}$, where $t_{\text{decay}}$ is the time it takes the black hole to decay. For charged black holes $t_{\text{decay}}$ must be calculated numerically. In Figure 6, these constraints largely follow curves of constant $t_{\text{cool}}+t_{\text{decay}}$, as the decay bounds usually depend on the timing of energy injections. The constraints disappear when $t_{\text{cool}}$ is larger than the age of the Universe, as these black holes have yet to form. The constraints consistently weaken for the highest values of $m_{X}$. This is because the black holes formed from heavier dark electrons tend to be closer to their extremal mass, and will emit a smaller fraction of their total mass ($\sim 10\%$) before becoming extremal. We give more detailed descriptions of how we calculated these constraints in Appendix B.
Black holes that collapse quickly and take longer than the lifetime of the universe to decay happen to be in a mass range where the charge of the black hole does not alter the observed radiation today. These are subject to the usual radiation constraints for black holes with lifetimes older than the Universe [74].
While we have estimated the constraints on charged late forming black holes for reference, these constraints generally rely on the entire emission history of the black holes. One would have to consider these black holes’ unique evolution history (rather than taking averages), to get precise constraints. Recently, a numerical analysis of constraints on the PBH abundance from BBN was studied in Ref. [75], which placed the strongest bounds on the PBH mass between $10^{8}$ g and $10^{9}$ g. An interesting extension would be to study how these bounds differ in the case of the delayed black holes presented in this work. We leave these more careful treatments to future work.
Furthermore, a particularly interesting and novel observable feature of this scenario is the late-time flashes from rapid evaporation of light PBHs [76, 77] which can be searched for in gamma ray detectors. While the charges on these black holes will prevent them from decaying to arbitrarily low masses (high temperatures), they can still get hot enough to probe energies inaccessible to colliders on Earth. Late-time forming black holes with light masses have been suggested as a possible solution to antiproton, gamma-ray, and Helium excesses from the galactic center [78, 79].
@@
Finally, in addition to the GW emission from the exotic cosmology, binary mergers of both the compact objects and black holes produced in this scenario would give rise to GW emissions [80, 81, 82, 83, 84]. Furthermore, the evaporation of PBHs can also affect the background of stochastic GWs [32, 85]. We leave these considerations for future studies. We comment that there are also dynamical constraints on compact objects, such as the disruption of dwarf galaxies via dynamical heating [86, 87] and wide binaries [88, 89]. However, these only become relevant at compact object masses larger than $\sim 10^{34}$ g, and are weaker than microlensing constraints until $\sim 10^{36}$ g, which is larger than any of the final object masses we have considered in this work. Other dynamical effects, such as the accumulation of compact objects in the galactic nucleus via dynamical friction [90] can also place constraints. Once again however, these become relevant at compact object masses higher than the ones in this work.